### v2

@@ 5.1 Setup
This generalized bounded noise [^4] The noise is bounded when the gradient is bounded but can become unbounded if the gradient explodes. assumption is inspired by the affine variance condition initially proposed by Bottou et al. (2018), where the noise level depends on the norm of the true gradient and is captured via an affine function. Assumption 4 encompasses the uniformly bounded noise (Zhang et al., 2020b) and affine noise (Liu et al., 2023b; Hong and Lin, 2024) condition with $\sigma(\alpha)\equiv\sigma$ and $\sigma(\alpha)\equiv\sigma_{0}+\sigma_{1}\alpha$, respectively. Moreover, we underline that Assumption 4 is never stronger than the commonly assumed finite moment or sub-Gaussian noise condition, with thorough discussions deferred to Section H.1.
5.2 Theoretical Guarantees
Extension from deterministic optimization to SCO is highly non-trivial, for the sake of the following two quantities that need to be controlled.
Suboptimality gap
Due to the inherent stochasticity, the sequence $\left\{f(\mathbf{x}_{t})-f^{*}\right\}_{t\in\mathbb{N}}$ may become unbounded. Recall from deterministic optimization, we directly link suboptimality gaps to the last-iterate convergence. However, stochastic gradient descent (SGD) (Robbins and Monro, 1951) primarily focuses on the average-iterate, and therefore, such a link does not hold. A fruitful line of work (Shamir and Zhang, 2013; Harvey et al., 2019; Orabona, 2020; Jain et al., 2021) explore the last-iterate convergence of SGD but often remain confined to Euclidean settings, compact domains, or uniformly bounded noise. Although Liu and Zhou (2024) overcome some of these limitations, their reliance on sub-Gaussian noise makes the analysis incompatible with generalized smoothness. Our solution is motivated by Orabona (2020), who elegantly converts average-iterate convergence of SGD to last-iterate in-expectation convergence for non-smooth objectives within Euclidean domains. Through a careful analysis of martingales and a refinement of their techniques, we extend the framework to $\ell*$-smooth objectives in non-Euclidean settings. Leveraging this conversion framework, we prove that suboptimality gaps are bounded for all iterations simultaneously with high probability, indicating that the gradients are also bounded for all $t\in\mathbb{N}$. Such “anytime” characteristic is vital for our analysis, since Lemma 3 needs to be applied to every round of SMD.
Dual norm of the stochastic gradient
In the context of SCO, we still need to exploit the local smoothness property (Lemma 3), which requires the distance between consecutive iterates to be bounded. In the deterministic case, we have $\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|\leq\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}$, and by Lemma 7, the boundedness of $\left\|\nabla f(\mathbf{x}_{t})\right\|_{*}$ is inferred from finite suboptimality gaps. In contrast, for the stochastic case, it holds that $\left\|\mathbf{x}_{t+1}-\mathbf{x}_{t}\right\|\leq\left\|\mathbf{g}_{t}\right\|_{*}$, where the challenge lies in bounding $\left\|\mathbf{g}_{t}\right\|_{*}$. Clearly, the self-bounding property in Lemma 7 fails to bridge the stochastic gradient and the suboptimality gap. To address this, we leverage the noise model in Assumption 4 to connect the noise level with the dual norm of the true gradient. As the suboptimality gap is bounded with high probability, the boundedness of the true gradient follows directly, implying that the noise is almost surely bounded via $\sigma(\cdot)$. Finally, we apply a union bound over these events to establish bounds on the stochastic gradient.
Based on the above reasoning, our analysis framework significantly differs from that of Li et al. (2023a, Section 5.2), who utilize a stopping-time technique (Williams, 1991) to show that the two key quantities are bounded before the stopping time $T$ with high probability. However, their method has two limitations: (i) it requires prior knowledge of the total number of iterations $T$; (ii) it lacks a convergence guarantee for the outputs generated before the final round. Moreover, their approach is ill-suited for SMD, as it heavily relies on the Euclidean inner product such that $\langle\mathbf{x},\mathbf{x}\rangle=\left\|\mathbf{x}\right\|_{2}^{2}$. Our main result is presented in Theorem 12, whose proof is outlined in Section 6.5.
Theorem 12.
Under Assumptions 1, 2, 3 and 4, set
$0<\eta_{t}\leq\min\left\{\frac{\min\left\{1,\left\|\nabla f(\mathbf{x}_{t-1})\right\|_{*}/\sigma(\left\|\nabla f(\mathbf{x}_{t-1})\right\|_{*})\right\}}{2\ell(2\left\|\nabla f(\mathbf{x}_{t-1})\right\|_{*})},\frac{D}{\sigma(\left\|\nabla f(\mathbf{x}_{t-1})\right\|_{*})\sqrt{t}}\right\}.$ (22)
Remark 5.1.
Theorem 12 achieves the classic $\widetilde{O}\left(1/\sqrt{t}\right)$ anytime convergence rate (Zhang et al., 2024a, Theorem 3.6) for SCO. To the best of our knowledge, the extra-logarithmic factors are unimprovable in this setting (Zhang et al., 2024b, Theorems 2 and 8). Note that existing results rely on a deterministic learning rate schedule proportional to $1/\sqrt{t}$. However, this strategy is too coarse to capture the local features of the objective under $\ell*$-smoothness. Consequently, we incorporate local curvature information into the learning rates, akin to the approach in Xie et al. (2024b), who consider online convex optimization under $\ell$-smoothness.
6 Analysis
We present the proofs of the main theorems in Sections 4 and 5, where the lemmas involved can be found in Appendices D, E, F, G and H.2.
6.1 Mirror Descent
Proof 6.1.

### v3

@@ Appendix C The Gain of Algorithmic Adaptivity
C.1 Theoretical Justifications: Non-constant Link Functions
C.1.1 Unnormalized Softmax Logits
Consider the following logistic kernel function:
[…]
We focus on the unbounded domain like $\mathbb{R}^{n}$, which is often the case in reality. For bounded domains (satisfying Assumption J.2) considered in this paper, the discussions are deferred to Section C.1.3. We will first show that the Hessian $\nabla^{2}f(\mathbf{x})$ is unbounded unless $\mathbf{w}=0$, indicating $f$ is not $L$-smooth for any constant $L$. Then, we demonstrate that generalized smoothness can effectively handle this problem. Finally, we give characterizations of $\ell*$-smoothness and $\ell$-smoothness, highlighting the edge of our formulation.
The gradient and Hessian of $f$ are given by
Then we have
$\ell$-smoothness
We have
Thus,
which implies $f\in\mathcal{F}_{\widehat{\ell}}(\left\|\cdot\right\|_{2})$ with $\widehat{\ell}(\alpha)=\left\|\mathbf{w}\right\|_{2}\alpha$. Obviously, $(L_{0},L_{1})$-smoothness can model the curvature of this function more appropriately.
$\ell*$-smoothness
Thus, $f\in\mathcal{F}_{\widetilde{\ell}}(\left\|\cdot\right\|_{1})$ with $\widetilde{\ell}(\alpha)=\left\|\mathbf{w}\right\|_{\infty}\alpha$.
Comparison
Define the ratio between $\ell*$-smoothness and $\ell$-smoothness as
From above, we can assert that our formulation is strictly better than that in Li et al. (2023a). Taking $\mathbf{w}=\mathbf{1}_{n}$ satisfies the left side of the inequality. In this case, $f\in\mathcal{F}_{\widehat{\ell}}(\left\|\cdot\right\|_{2})$ with $\widehat{\ell}(\alpha)=\sqrt{n}\alpha$ and $f\in\mathcal{F}_{\widetilde{\ell}}(\left\|\cdot\right\|_{1})$ with $\widetilde{\ell}(\alpha)=\alpha$. Clearly, our definition surpasses $\ell$-smoothness by a dimension-dependent factor.
Even more practical case
This time, we consider $\mathbf{w}\in\mathbb{R}^{n}$ as a random variable uniformly distributed on a unit sphere. Since the slopes of the generalized smooth functions $\widehat{\ell}$$\widetilde{\ell}$ are now randomized, we measure the smoothness parameter in expectation. We clearly have $\left\|\mathbf{w}\right\|_{2}=1$ by definition. As for $\widetilde{\ell}$, it is well-known (Schechtman and Zinn, 2000; Boucheron et al., 2013) that
Hence, the (expected) ratio becomes
Mathematically, we establish the bound $\widetilde{\ell}/\widehat{\ell}\preceq\sqrt{\log(n)/n}$ in expectation, demonstrating the quantitative advantage of $\ell*$-smoothness. This benefit becomes particularly pronounced when handling high-dimensional optimization problems, mirroring practical scenarios where our adaptive approach offers meaningful improvements over classical smoothness analysis.
C.1.2 Logistic Regression
Consider the loss function of logistic regression:
[…]
whose gradient and Hessian are given by
Thus we have
implying that $f$ is $\left|C\right|\left\|\mathbf{w}\right\|_{2}^{2}/4$-smooth. This smoothness constant can become very large if $\left|C\right|\cdot\left\|\mathbf{w}\right\|_{2}\gg 1$. Therefore, it’s meaningful to seek an alternative. Let’s consider the generalized smoothness notion as in Section C.1.1.
$\ell$-smoothness
We have
Thus,
Since $\left\|\nabla^{2}f(\mathbf{x})\right\|_{2}\leq\left\|\mathbf{w}\right\|_{2}\cdot\left\|\nabla f(\mathbf{x})\right\|_{2},\forall\mathbf{x}\in\mathbb{R}^{n}$, we conclude that $f\in\mathcal{F}_{\widehat{\ell}}(\left\|\cdot\right\|_{2})$ with $\widehat{\ell}(\alpha)=\left\|\mathbf{w}\right\|_{2}\alpha$ (Chen et al., 2023b). Clearly, when $\left\|\mathbf{w}\right\|_{2}\gg 1$, generalized smooth is more appropriate because it improves upon the standard smooth by a factor of $\left\|\mathbf{w}\right\|_{2}$.
$\ell*$-smoothness
We follow a similar procedure as above:
and compute
Therefore, we conclude that $f\in\mathcal{F}_{\widetilde{\ell}}(\left\|\cdot\right\|_{1})$ with $\widetilde{\ell}(\alpha)=\left\|\mathbf{w}\right\|_{\infty}\alpha$.
Comparison
Similarly, we can define the ratio function $\phi(\mathbf{w})$ as in Section C.1.1, and the same conclusions can be drawn: our definition surpasses $\ell$-smoothness by a dimension-dependent factor of $1/\sqrt{n}$.
Logistic regression on a high-dimensional sphere
Similar to Section C.1.1, we now consider logistic regression on a high-dimensional sphere, with radius $R\gg 1$. In this case, $\left\|\mathbf{w}\right\|_{2}=R\gg 1$, and $\mathbf{w}$ is uniformly distributed on the $R$-radius sphere. Note that $\mathbf{w}/R$ can be seen as a random variable uniformly distributed on the unit sphere, then we we also have $\phi_{\mathbb{E}}(\mathbf{w}):=\frac{\mathbb{E}[\left\|\mathbf{w}\right\|_{\infty}]}{\mathbb{E}[\left\|\mathbf{w}\right\|_{2}]}=\mathbb{E}[\left\|\mathbf{w}\right\|_{\infty}]\preceq\sqrt{\frac{\log n}{n}}$. Again, we verify the effectiveness and benefits of our $\ell*$-smoothness in this practical scenario.
C.1.3 The cases for bounded domains
The examples and justification in Sections C.1.1 and C.1.2 targets unbounded domains like $\mathbb{R}^{n}$. Under Assumption J.2, the previous arguments are still valid and we emphasize that the advantage of $\ell*$-smoothness does not stem from unboundedness of the domain.
Unnormalized Softmax Logits
For $f(\mathbf{x})=C\exp(\mathbf{w}^{\top}\mathbf{x}+b),\mathbf{w},\mathbf{x}\in Y:=\{\mathbf{y}\in\mathbb{R}^{n}|\|\mathbf{y}\|_{2}\leq R\}$. Then the standard smoothness constant is $L=\max_{\mathbf{x}}\|\nabla^{2}f(\mathbf{x})\|_{2}\leq\left|C\right|\exp(R^{2}+b)R^{2}$. Unfortunately, it exhibits an exponential dependence on $R$, making traditional smoothness impractical. For instance, when $\left|C\right|=b=R=5$, we get $L>10^{15}$, which is clearly meaningless and necessitates a better formulation. One can verify that we still have $f\in\mathcal{F}_{\widehat{\ell}}(\left\|\cdot\right\|_{2})$ with $\widehat{\ell}(\alpha)=\|\mathbf{w}\|_{2}\alpha$ and $f\in\mathcal{F}_{\widetilde{\ell}}(\left\|\cdot\right\|_{1})$ with $\widetilde{\ell}(\alpha)=\|\mathbf{w}\|_{\infty}\alpha$.
Logistic Regression
Main takeaways
C.2 Theoretical Justifications: Constant Link Functions
Consider the function $h(\mathbf{x})=\mathbf{x}^{\top}(\mathbf{1}_{n}-\mathbf{e}_{1})(\mathbf{1}_{n}-\mathbf{e}_{1})^{\top}\mathbf{x}/2,\mathbf{x}\in\Delta_{n}$, where $\Delta_{n}=\{\mathbf{x}\geq\mathbf{0},\mathbf{1}_{n}^{\top}\mathbf{x}=1|\mathbf{x}\in\mathbb{R}^{n}\}$ is the (n1)(n-1)-dimensional simplex. We have the following proposition.
Proposition C.1.
(i) $h\in\mathcal{F}_{\widehat{\ell}}(\left\|\cdot\right\|_{2})$ with $\widehat{\ell}(\cdot)\equiv n-1$; (ii) $h\in\mathcal{F}_{\widetilde{\ell}}(\left\|\cdot\right\|_{1})$ with $\widetilde{\ell}(\cdot)\equiv 1$.
Proof.
$\nabla h(\mathbf{x})=A\mathbf{x},\quad\nabla^{2}h(\mathbf{x})=A.$ (21)
[…]
We proceed by computing
where we make use of $\left\|\mathbf{1}_{n}-\mathbf{e}_{1}\right\|_{\infty}=1$. Therefore, we conclude that $\widetilde{\ell}(\cdot)\equiv 1$. ∎
Under $\ell$-smoothness, the function $h$ inevitably has a smoothness parameter that scales linearly to the dimension $n$. Whereas, under $\ell*$-smoothness, the dependence of the dimension no longer exists. The additional factor n1n-1 will inevitably incur a dimension-dependent convergence rate, which will be stated explicitly in Section C.4.
C.3 Empirical Justifications
(b) Comparison of $\mathcal{F}_{\widehat{\ell}}(\left\|\cdot\right\|_{2})$ and $h\in\mathcal{F}_{\widetilde{\ell}}(\left\|\cdot\right\|_{1})$ generalized smooth functions class for different values of $n$.
We focus on the following example function:
[…]
$\widehat{\ell}(\alpha)=\widehat{L}_{0}+\widehat{L}_{1}\alpha,\quad\widetilde{\ell}(\alpha)=\widetilde{L}_{0}+\widetilde{L}_{1}\alpha.$ (27)
We find this characterization reasonable based on observations from Figure 5(b): when $n>20$, the scattered points align into a straight line. To this end, we fit the values of $\widehat{L}_{0}$, $\widehat{L}_{1}$, $\widetilde{L}_{0}$, and $\widetilde{L}_{1}$ using $500$ randomly sampled points from the simplex $\Delta_{n}$.
Based on the modeling in (27), we can see that the slopes $\widehat{L}_{1}$ and $\widetilde{L}_{1}$ determine the magnitude of smoothness along the optimization trajectory (e.g., for some constant $0<G<\infty$, $\widehat{\ell}(G)=\widehat{L}_{0}+\widehat{L}_{1}G$), to a great extent. To clarify, note that the definition of $G$ in (5) has a closed-form solution:
[…]
C.4 Comparison on the Convergence Rates
[…]
$\frac{\widehat{\ell}(\left\|\nabla f(\mathbf{x}_{0})\right\|_{2})\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|_{2}^{2}}{2T}\leq\frac{\widehat{\ell}(\left\|\nabla f(\mathbf{x}_{0})\right\|_{2})}{2T}\triangleq R_{1}(T),$ (30)
where we use $\left\|\mathbf{x}_{0}-\mathbf{x}_{*}\right\|_{2}^{2}\leq\sup_{\mathbf{x},\mathbf{y}\in\Delta_{n}}\left\|\mathbf{x}-\mathbf{y}\right\|_{2}^{2}=1$. For mirror descent, as established in Theorem 3.5, the convergence rate is given by
$\frac{\widetilde{\ell}(G)B(\mathbf{x}_{*},\mathbf{x}_{0})}{T}\leq\frac{\widetilde{\ell}(G)\ln n}{T}\triangleq R_{2}(T),$ (31)
[…]
underlining the advantage of mirror descent over gradient descent by reducing the dependency on the dimensionality factor. Obviously, the justifications for non-constant link functions in Section C.1 also follow from the above discussions, which we omit for simplicity.
Appendix D Properties of Generalized Smooth Function Classes
In this section, we demonstrate some basic properties of the function class $\mathcal{F}_{\ell}(\left\|\cdot\right\|)$ as well as $\mathcal{F}_{\ell,r}(\left\|\cdot\right\|)$. Equation 34 is a generalized version of Lemma A.4 of Li et al. (2023a), which only supports $f\in\mathcal{F}_{\ell}(\left\|\cdot\right\|_{2})$. Since the original derivations did not utilize the inherent property of the Euclidean norm, i.e., $\langle\cdot,\cdot\rangle=\left\|\cdot\right\|_{2}^{2}$, we can adapt its proof to accommodate an arbitrary norm $\left\|\cdot\right\|$.
Lemma D.1.
Proof.
First, define $g(s):=\left\|\nabla f(\mathbf{x}(s))\right\|_{*}$ for $0\leq s\leq 1$. Since $\mathbf{x}(s)\in\mathcal{X}$, $\mathbf{g}(s)$ is differentiable almost everywhere by Definition 2.1. Hence, the following relation holds almost everywhere over $(0,1)$:
where the first inequality uses the triangle inequality and the last inequality uses Definition 2.1. Next, we define $h(u):=\ell(u)\left\|\widetilde{\mathbf{x}}-\mathbf{x}\right\|$ and $k(u):=\int_{0}^{u}1/h(v)\mathrm{d}v$. We immediately conclude that $g^{\prime}(s)\leq h(g(s))$ holds almost everywhere over $(0,1)$, which satisfies the condition of a generalized version of Grönwall’s inequality (Li et al., 2023a, Lemma A.3). Hence, we have
Based on this result, we can derive
Maximizing over $0\leq\tilde{s}\leq 1$ yields the desired result. ∎
In the following, we provide the omitted proof of Proposition 2.6.
Proof of Proposition 2.6.
The proof is very similar to that of Li et al. (2023a, Proposition 3.2), except that the derivations based on $\left\|\cdot\right\|_{2}$ should be replaced by an arbitrary norm $\left\|\cdot\right\|$.
Given an $\mathbf{x}\in\mathcal{X}$ where $\nabla^{2}f(\mathbf{x})$ exists, according to Definition 2.4, the following relation holds
  1. 1.
  2. 2.
[…]
By the definition of $s_{b}$ and Assumption 2.5, we have $\lim_{s\to s_{b}}f(\mathbf{x}(s))=\infty$ and $\mathbf{x}(s)\in\mathcal{X}$ for $0\leq s<s_{b}$. Then we have
by definition. This contradicts $\lim_{s\to s_{b}}f(\mathbf{x}(s))=\infty$. Thus we have $\widetilde{\mathbf{x}}\in\mathcal{X}$ and consequently, $\mathcal{B}(\mathbf{x},\widetilde{r}(\left\|\nabla f(\mathbf{x})\right\|_{*}))\subseteq\mathcal{X}$. Next, we prove (2) in Definition 2.4. For any $\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{B}(\mathbf{x},\widetilde{r}(\left\|\nabla f(\mathbf{x})\right\|_{*}))$, define $\mathbf{x}_{1,2}(s):=s\mathbf{x}_{1}+(1-s)\mathbf{x}_{2}$ for $0\leq s\leq 1$. Since $\mathbf{x}_{1,2}(s)\in\mathcal{B}(\mathbf{x},\widetilde{r}(\left\|\nabla f(\mathbf{x})\right\|_{*}))$, we obtain
where the first inequality uses Definition 2.1 and the last inequality uses Equation 34 and the non-decreasing property of $\ell$. ∎
We present the following effective or local smoothness property derived under generalized smoothness. Lemma 2.7 is extremely useful as it closes the gap between generalized smoothness and classical smoothness.
Proof of Lemma 2.7.
To prove this lemma, we adapt the existing result for $f\in\mathcal{F}_{\ell,r}(\left\|\cdot\right\|_{2})$ (Li et al., 2023a, Lemma 3.3). Specifically, we establish three analogous statements for $f\in\mathcal{F}_{\ell,r}(\left\|\cdot\right\|)$ and then leverage the equivalence between generalized smooth function classes to conclude that these statements also hold for $f\in\mathcal{F}_{\ell}(\left\|\cdot\right\|)$. By Definition 2.4, we have $\ell(\left\|\nabla f(\mathbf{x})\right\|_{*})\leq\ell(G)$ and $r(G)\leq r(\left\|\nabla f(\mathbf{x})\right\|_{*})$. Hence, it follows that (i) $\mathcal{B}(\mathbf{x},r(G))\subseteq\mathcal{B}(\mathbf{x},r(\left\|\nabla f(\mathbf{x})\right\|_{*}))\subseteq\mathcal{X}$; and (ii) $\forall\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{B}(\mathbf{x},r(G))$:
At last, we apply the equivalence of generalized smooth function classes to show that the same derivations also hold for $f\in\mathcal{F}_{\ell}(\left\|\cdot\right\|)$. In this case, it suffices to set $L=\ell(2G)$ and $r(G)=G/L$ according to Proposition 2.6. Then the statements (i), (ii), and (iii) can be easily transformed into the ones in Lemma 2.7, which concludes the proof. ∎
Lemma D.2.
Under Assumption 3.3, let $\left\|\nabla f(\mathbf{x})\right\|_{*}\leq G\in\mathbb{R}_{+}$ for any given $\mathbf{x}\in\mathcal{X}$. If $0<\eta\leq\gamma/\ell(2G)$ for some $\gamma\in(0,1)$, then for any $\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{B}(\mathbf{x},G/\ell(2G))$ and any $\mathbf{x}_{3},\mathbf{x}_{4}\in\mathcal{X}$,
holds for any $\beta\in\mathbb{R}_{++}$.
Proof.
Since $\left\|\nabla f(\mathbf{x})\right\|_{*}\leq G$ and $\mathbf{x}_{1},\mathbf{x}_{2}\in\mathcal{B}(\mathbf{x},G/\ell(2G))$, by Lemma 2.7, we have
According to Hölder’s Inequality, we have
Combining these two inequalities with the condition that $0<\eta\leq\gamma/\ell(2G)$ completes the proof. ∎
Lastly, we provide the omitted proof of Lemma 3.4 in Section 3.2Lemma 3.4 is also a generalized version of the reversed Polyak-Łojasiewicz Inequality in Li et al. (2023a, Lemma 3.5). Since their derivations rely on the inherent properties of the Euclidean norm, we need to take a different approach to prove this lemma.
Proof of Lemma 3.4.
For any $\mathbf{x}\in\mathcal{X}$ and any $\mathbf{h}$ satisfying $\left\|\mathbf{h}\right\|\leq\left\|\nabla f(\mathbf{x})\right\|_{*}/\ell(2\left\|\nabla f(\mathbf{x})\right\|_{*})$, by Lemma 2.7, the following relation holds:
Lemma D.3.
Under Assumption 3.3, for a given $\mathbf{x}\in\mathcal{X}$ satisfying $f(\mathbf{x})-f^{*}\leq F$, we have $\left\|\nabla f(\mathbf{x})\right\|_{*}\leq G<\infty$ and $G^{2}=2\ell(2G)F$, where $G:=\sup\left\{\alpha\in\mathbb{R}_{+}|\alpha^{2}\leq 2F\ell(2\alpha)\right\}$.
Proof.
Based on Lemma 3.4, the proof is analogous to that of Li et al. (2023a, Corollary 3.6), which involves merely an $f\in\mathcal{F}_{\ell}(\left\|\cdot\right\|_{2})$. Since we have a sub-quadratic $\ell$, we clearly have $\lim_{\alpha\to\infty}\frac{\alpha^{2}}{2\ell\left(2\alpha\right)}=\infty$. So for $F\in\mathbb{R}_{+},\exists\alpha_{F}\in\mathbb{R}_{++}$, such that $\forall\alpha>\alpha_{F},\frac{\alpha^{2}}{2\ell\left(2\alpha\right)}>F$. Therefore, if $\frac{\alpha^{2}}{2\ell\left(2\alpha\right)}\leq F$ holds for some $\alpha$, then it must be that $\alpha\leq\alpha_{F}$. Hence, our construction of $G$ implies that $G\leq\alpha_{F}<\infty$. We conclude that $\left\|\nabla f(\mathbf{x})\right\|_{*}\leq G$ by Lemma 3.4. Lastly, $G^{2}=2\ell(2G)F$ is directly implied by the compactness of the set $\left\{\alpha\in\mathbb{R}_{+}|\alpha^{2}\leq 2F\ell(2\alpha)\right\}$. ∎
Appendix E Analysis for Mirror Descent
We present the following standard lemma for mirror descent, which can be found in Bubeck and others (2015) and Lan (2020).
Lemma E.1.
Proof.
Take $\mathbf{x}=\mathbf{x}_{0}$ and rearrange it,
If $\left\|\mathbf{x}_{0}-\mathbf{x}^{+}\right\|=0$Lemma E.1 is naturally true. Otherwise, dividing both sides of the inequality by $\left\|\mathbf{x}_{0}-\mathbf{x}^{+}\right\|$ yields the result. ∎
Lemma E.2.
Under Assumptions 2.5 and 3.3, for any $\mathbf{x}_{0}\in\mathcal{X}$ satisfying $f(\mathbf{x}_{0})-f^{*}\leq F$, define $G:=\sup\left\{\alpha\in\mathbb{R}_{+}|\alpha^{2}\leq 2F\ell(2\alpha)\right\}$ and $\mathbf{x}^{+}=\mathcal{P}_{\mathbf{x}_{0}}(\eta\nabla f(\mathbf{x}_{0}))$. If $0<\eta\leq 1/\ell(2G)$, we have $f(\mathbf{x}^{+})\leq f(\mathbf{x}_{0})$ and $\left\|\nabla f(\mathbf{x}^{+})\right\|_{*}\leq G<\infty$.
Proof.
Under Assumption 2.5, we know that $f^{*}>$-$\infty$, which gives $f(\mathbf{x}_{0})-f^{*}<\infty$. By Lemma D.3, we immediately deduce that $\left\|\nabla f(\mathbf{x}_{0})\right\|_{*}\leq G<\infty$. Then we invoke Lemma E.1 to control the distance between two iterates:
where we use $\eta\leq 1/\ell(2G)$. The conditions of Lemma 2.7 are now satisfied, so we use the effective smoothness property to obtain
Multiplying by $\eta$ and adding this to (57), we have
$\leq$ $\eta f(\mathbf{x}_{0}),$
$f(\mathbf{x}^{+})-f^{*}\leq f(\mathbf{x}_{0})-f^{*}.$ (61)
Invoking Lemma D.3 again finishes the proof. ∎
Proof of Theorem 3.5.